# Optical Properties, Quasinormal Modes and Greybody factors of deformed AdS-Schwarzschild black holes

### v1

@@ Abstract
We investigate the temperature, photon and shadow radii, quasinormal modes (QNMs), greybody factors, and emission rates of deformed AdS black holes, focusing on the effects of the deformation parameter $\alpha$ and control parameter $\beta$. Increasing $\alpha$ enhances the oscillation frequency and damping rate of gravitational waves, while $\beta$ shows non-linear behaviour. Electromagnetic perturbations exhibit similar trends, though with lower frequencies and damping rates. Greybody factors are mainly influenced by multipole moment $l$ and $\alpha$, with $\beta$ having a more subtle effect. These findings provide insights into black hole dynamics, mergers, and gravitational wave emissions.
WKB method; Black hole; Quasinormal modes, Greybody factors
@@ I Introduction
Black hole thermodynamics began with a groundbreaking insight by J. D. Bekenstein, who, in the early 1970s, proposed that black holes should possess entropy. His work established a direct relationship between the area of a black hole’s event horizon and its entropy, indicating that the surface area of the horizon could be a measure of the disorder or information content of a black hole Bekenstein:1972tm . This idea transformed the way scientists viewed black holes, not merely as objects that consume matter and energy, but as thermodynamic entities that obey laws analogous to those of classical thermodynamics.
@@
In recent years, research has continued to explore black hole thermodynamics across a variety of theoretical frameworks, including modified theories of gravity such as string theory and loop quantum gravity Ashtekar:2004eh ; Dvali:2011aa . These modified theories often lead to generalizations or deformations of the traditional thermodynamic laws that apply to black holes. For example, when incorporating the generalized uncertainty principle, a modification of the Heisenberg uncertainty principle to account for quantum gravitational effects, the entropy and temperature of black holes are modified, leading to potentially observable consequences Amelino-Camelia:2000stu ; Ali:2009zq ; Yang:2023nnk . These generalized uncertainty principles are particularly relevant in the context of small black holes or black holes that form in high-energy regimes, such as those potentially produced in particle accelerators.
@@
For example, research on rotating Kerr-AdS black holes using holographic techniques has provided insights into their phase transitions and thermodynamic properties Gong:2023ywu . Similarly, accelerating black holes and those in lower-dimensional spacetimes offer simplified contexts for studying thermodynamic stability and transitions, contributing to the broader understanding of black hole microstructures Sokoliuk:2023pby ; Singh:2023hit ; Blagojevic:2023kqj ; Xiao:2023lap ; Wu:2022whe . Such studies are crucial for understanding the nature of black hole entropy and the information paradox, one of the most significant unresolved issues in theoretical physics.
Phase transitions and critical phenomena have been extensively investigated in black hole thermodynamics Gunasekaran:2012dq ; Cvetic:1999ne ; Chamblin:1999tk . Various types of black holes in different modified gravity theories have also been found to exhibit a wide range of phase transitions Wei:2012ui ; Altamirano:2014tva .
@@
In this paper, we adopt the metric signature (,+,+,+)(-,+,+,+)( - , + , + , + ) and use geometrical units where $c=G=\hbar=1$. After this brief introduction, the paper is structured as follows:
In Section II, we provide a concise overview of the deformed black hole solution. Section III examines the Hawking temperature of the black hole and explores the remnant radius and mass. In Section IV, we analyze the photon sphere radius based on the motion of light and derive the black hole’s shadow radius. Section V investigates the quasinormal modes (QNMs) of the black hole spacetime, while Section VI focuses on the greybody factors. The emission rate of the black hole is discussed in Section VII. Finally, in Section VIII, we conclude with a brief summary and discussion.
@@ II Deformed Black hole solution
To construct the deformed black hole solution, we consider a static, spherically symmetric spacetime with the line element
@@
where $\alpha$ and $\beta$ are deformation parameters. The parameter $\alpha$ controls the strength of the deformation, while $\beta$ ensures the avoidance of central singularities at $r=0$. This choice of energy density function satisfies the necessary energy conditions and asymptotically decays as $r\to\infty$.
@@ III Temperature
At horizon $\mathcal{F}(r)=0\big{|}_{r=r_{h}}$, thus one can find the mass of the black hole as a function of the horizon radius.
@@
Hawking temperature is obtained from $\frac{1}{4\pi}\partial\mathcal{F}(r)/\partial r|_{r=r_{h}}$. For the metric of the form 6, by substituting mass from Eq. (7), the temperature would be found as
$T_{H}=\frac{\beta^{4}+r^{4}+4\beta r^{3}-\alpha r^{2}+6\beta^{2}r^{2}-\Lambda r^{2}(\beta+r)^{4}+4\beta^{3}r}{4\pi r(\beta+r)^{4}}$ (8)
@@
By substituting remnant mass from Eq. (9), by $r_{h}$ in Eq. (7), remnant mass of the black hole can be find as
@@ IV Photon and Shadow Radii
where Lagrangian is obtained as Ahmed
@@
Thus one can define an effective potential which is formulated as
@@
To ensure a circular orbit, two constraints are applied which are as follow
@@
Thus one can find that the following constraints should be satisfied
$E/L$ =±$=\pm\sqrt{\mathcal{F}(r)/r^{2}},$ (17)
@@
$r_{sh}=r_{ph}\frac{\sqrt{f(r_{o})}}{\sqrt{f(r_{ph})}}$ (19)
that $r_{o}$ indicates the observer radius. Shadow radius is shown in Figs. 3 and 4. Table 1, represents the permissible range of parameters based on the observation of Sgr A for $1\sigma$ and $2\sigma$ regions.
@@
Figure 3: Variation of shadow radius with parameter $\alpha$ for $r_{o}=10M$. In the left panel $\beta=0.8$ and in the right panel $\Lambda=-0.002$ are considered.
@@
As evident in Figs. 3 to 5, a reduction in the parameters $\alpha$, and $\Lambda$, while the rest of the parameters remain constant, results in a decrease in the value of the shadow radius.
The allowed range of parameters $\alpha$ and $\Lambda$, where other parameters are constant, in which shadow radius lies in the $1\sigma$ and $2\sigma$ regions, due to the observations of Sgr A is represented in table 1. Also, as shown in Fig. 5, by setting $\alpha=0.2$ and for the given $\Lambda$s, the value of shadow radius lies in the $1\sigma$ region of Sgr A. By setting $\Lambda=-0.002$, for the case $\alpha=0.50$, for $\beta>0.83$, shadow radius is permissible in $1\sigma$ region due to the observations of Sgr A.
@@
$\beta=0.900$--- - - ---- - - $0.720$--- - - $\Lambda=-0.002$--- - - ---- - - $0.493$--- - -
$\beta=0.030$--- - - ---- - - $0.508$ $0.813$ $\Lambda=-0.006$--- - - ---- - - $0.959$--- - -
$\beta=0.010$--- - - ---- - - $0.415$ $0.653$ $\Lambda=-0.010$--- - - $0.156$--- - - ---- - -
@@
$\alpha=0.900$ $-0.041$ $-0.030$ $-0.014$ $-0.007$ $\beta=0.900$ $-0.022$ $-0.014$ $-0.003$--- - -
$\alpha=0.500$ $-0.025$ $-0.017$ $-0.005$--- - - $\beta=0.500$ $-0.024$ $-0.016$ $-0.004$--- - -
$\alpha=0.010$ $-0.015$ $-0.009$--- - - ---- - - $\beta=0.050$ $-0.028$ $-0.019$ $-0.006$ $-0.001$
@@ V Massless Scalar and Electromagnetic Perturbations in Black Hole Spacetimes
so that, now, $l$ becomes the multipole moment of the black hole’s QNMs. The frequency $\omega$ in the stationary Schrödinger equation (21) is defined via $\partial_{t}^{2}\psi_{sl}(t,r)=-\omega^{2}\psi_{sl}(t,r)$ and this definition is made possible by the diagonal and static nature of the metric in (4). The wave function $\psi(r_{*})_{s}$ stands for the waves associated with the scalar QNMs Chandrasekhar:1985kt ; lopez2020 .
@@
This effective potential $\mathcal{V}_{em}(r)$ of vector perturbations differs from the effective potential $\mathcal{V}_{s}(r)$ of scalar perturbations by the presence of the derivative of $g_{tt}$ in the latter.
@@
In Fig.s 6, 7 and 8, we have depicted the behaviours of the scalar and electromagnetic potentials with respect to the model parameters $l$, $\alpha$ and $\beta$ respectively. From Fig. 6, we can see that the potentials associated with both scalar and electromagnetic potentials behave similarly. However, for the scalar perturbation, the value of the peak of the potential is higher than that of the electromagnetic potential.
@@
Table 2: The QNMs for the massless scalar and electromagnetic perturbation with $n=0$, $M=1$, $\alpha=0.3$ and $\beta=0.2$ using the Padé averaged 6th order WKB approximation method.
  $l$    Scalar QNMs   Electromagnetic QNMs    $\vartriangle_{rms}(scalar)$    $\vartriangle_{rms}(em)$    $\Delta_{6}(scalar)$    $\Delta_{6}(em)$
$l=1$ $0.310177-0.100055i$ $0.263763-0.0953521i$ $7.68488\times 10^{-7}$ $7.14411\times 10^{-6}$ $3.09809\times 10^{-6}$ $7.38536\times 10^{-6}$
$l=2$ $0.511573-0.0993462i$ $0.484444-0.0977656i$ $2.14225\times 10^{-7}$ $3.75447\times 10^{-7}$ $2.10528\times 10^{-6}$ $5.9346\times 10^{-6}$
$l=3$ $0.714112-0.0991488i$ $0.694861-0.0983548i$ $4.21465\times 10^{-8}$ $5.35672\times 10^{-8}$ $4.10747\times 10^{-7}$ $1.12862\times 10^{-7}$
$l=4$ $0.917035-0.0990674i$ $0.902102-0.0985899i$ $1.06887\times 10^{-8}$ $1.23224\times 10^{-8}$ $1.17441\times 10^{-7}$ $3.25009\times 10^{-7}$
$l=5$ $1.12013-0.0990262i$ $1.10793-0.0987075i$ $3.43245\times 10^{-9}$ $3.78076\times 10^{-9}$ $1.2104\times 10^{-8}$ $1.19967\times 10^{-8}$
@@
We have listed QNMs for $n=0$ and for different $l$ values in Table 2 for both scalar and electromagnetic perturbations. These results were obtained for deformation parameters $\alpha=0.3$ and $\beta=0.2$, which influence the spacetime geometry and black hole properties. The table displays QNM frequencies, with the real part representing the oscillation frequency and the imaginary part corresponding to the damping rate of the modes. These modes were calculated for different multipole numbers $l$, ranging from $l=1$ to $l=5$.
@@
As expected, the real part of the QNMs increases with $l$, indicating higher oscillation frequencies for modes with larger angular momentum. The imaginary part, which controls the damping or decay rate, shows slight variation across the different values of $l$. Interestingly, scalar perturbations consistently exhibit slightly higher oscillation frequencies and more pronounced damping compared to electromagnetic perturbations. This is reflected in the larger imaginary components for scalar modes compared to their electromagnetic counterparts, highlighting the different behaviours of these two types of perturbations in deformed AdS black hole spacetimes.
@@
The table also presents the root mean square (rms) errors, $\Delta_{rms}$, for both scalar and electromagnetic perturbations. These errors quantify the precision of the WKB method in estimating the QNMs. It is observed that the RMS errors generally decrease with increasing $l$, indicating more accurate QNM calculations for higher multipole numbers. The WKB errors, $\Delta_{6}$, which is basically an error term associated with half of the absolute difference between QNMs calculated using 5th and 7th order WKB methods, similarly decrease with increasing $l$. This suggests that the WKB method becomes more reliable for larger values of $l$, where the approximation yields more precise QNM frequencies. Nevertheless, the electromagnetic perturbations exhibit slightly larger errors compared to scalar perturbations, although both remain within acceptable limits, particularly for $l\geq 2$.
@@
In Fig. 9, we have shown the variation of QNMs with respect to the model parameter $\alpha$ for scalar perturbations. On the first panel, we have shown the variations of the real part of QNMs i.e. the oscillation frequencies of ring-down GWs with respect to the model parameter $\alpha$. With an increase in the value of $\alpha$, oscillation frequency increases significantly. In the case of imaginary QNMs, we can see that an increase in the value of $\alpha$ increases the damping rate of ring-down GWs. In Fig. 10, we have depicted the variations of QNMs with respect to the model parameter $\beta$. On the first panel of Fig. 10, one can see that with an increase in the value of $\beta$, the real part of QNMs i.e. oscillation frequency of GWs decreases non-linearly. In the case of damping rate, we observe an interesting behaviour. Initially, the damping rate of the decay rate of ring-down GWs increases with an increase in the value of $\beta$ up to around $\beta=0.65.$ Beyond this value, the decay rate decreases with an increase in the model parameter $\beta$. In the cases of electromagnetic perturbations, we observe a similar scenario with slightly lower values of oscillation frequencies as shown in Fig.s 11 and 12.

### v2

# Optical, Dynamic and Topological Characteristics of Deformed Schwarzschild Black Holes

@@ Abstract
We investigate the temperature, photon and shadow radii, quasinormal modes (QNMs), time domain profiles, greybody factors, emission rates and topological characteristics of deformed black holes, focusing on the effects of the deformation parameter $\alpha$ and control parameter $\beta$. Increasing $\alpha$ enhances the oscillation frequency and damping rate of gravitational waves, while $\beta$ shows non-linear behaviour. Electromagnetic perturbations exhibit similar trends, though with lower frequencies and damping rates. Greybody factors are mainly influenced by multipole moment $l$ and $\alpha$, with $\beta$ having a more subtle effect. These findings provide insights into black hole dynamics, mergers, and gravitational wave emissions.
Emission Rate; Black hole; Quasinormal modes; Greybody factors
@@ I Introduction
Black hole thermodynamics began with a groundbreaking insight by J. D. Bekenstein, who, in the early 1970s, proposed that black holes should possess entropy Bekenstein:1972tm . His work established a direct relationship between the area of a black hole’s event horizon and its entropy, indicating that the surface area of the horizon could be a measure of the disorder or information content of a black hole Bekenstein:1972tm . This idea transformed the way scientists viewed black holes, not merely as objects that consume matter and energy, but as thermodynamic entities that obey laws analogous to those of classical thermodynamics.
@@
In recent years, researchers have continued to explore black hole thermodynamics across a variety of theoretical frameworks, including modified theories of gravity such as string theory and loop quantum gravity Ashtekar:2004eh ; Dvali:2011aa . These modified theories often lead to generalizations or deformations of the traditional thermodynamic laws that apply to black holes. For example, when incorporating the generalized uncertainty principle, a modification of the Heisenberg uncertainty principle to account for quantum gravitational effects, the entropy and temperature of black holes are modified, leading to potentially observable consequences Amelino-Camelia:2000stu ; Ali:2009zq ; Yang:2023nnk . These generalized uncertainty principles are particularly relevant in the context of small black holes or black holes that form in high-energy regimes, such as those potentially produced in particle accelerators.
@@
For example, research on rotating Kerr-AdS black holes using holographic techniques has provided insight into their phase transitions and thermodynamic properties Gong:2023ywu . Similarly, accelerating black holes and those in lower-dimensional spacetimes offer simplified contexts for studying thermodynamic stability and transitions, contributing to the broader understanding of black hole microstructures Sokoliuk:2023pby ; Singh:2023hit ; Blagojevic:2023kqj ; Xiao:2023lap ; Wu:2022whe . Such studies are crucial for understanding the nature of black hole entropy and the information paradox, one of the most significant unresolved issues in theoretical physics. Phase transitions and critical phenomena have been extensively investigated in black hole thermodynamics Gunasekaran:2012dq ; Cvetic:1999ne ; Chamblin:1999tk . Various types of black holes in different modified gravity theories have also been found to exhibit a wide range of phase transitions Wei:2012ui ; Altamirano:2014tva .
@@
In this paper, we adopt the metric signature (,+,+,+)(-,+,+,+) and use geometrical units where $c=G=\hbar=1$. After this brief introduction, the paper is structured as follows: In Section II, we provide a concise overview of the deformed black hole solution. Section III examines the Hawking temperature of the black hole and explores the remnant radius and mass. In Section IV, we analyze the photon sphere radius based on the motion of light and derive the black hole’s shadow radius. Section V investigates the quasinormal modes (QNMs) of the black hole spacetime, while Section VI focuses on the greybody factors. The emission rate of the black hole is discussed in Section VII. The topological characteristics of the black hole is studied in Section VIII. Finally, in Section IX, we conclude with a brief summary and discussion.
@@ II Deformed Black hole solution
To construct the deformed black hole solution, we consider the absence of matter field and a static, spherically symmetric spacetime with the line element
@@
where $\alpha$ and $\beta$ are deformation parameters. The parameter $\alpha$ controls the strength of the deformation, while $\beta$ ensures the avoidance of central singularities at $r=0$. This choice of energy density function satisfies the necessary energy conditions and asymptotically decays as $r\to\infty$. The specific form of $\mathcal{E}(r)$ in Eq. (5) is motivated by the requirement to construct a regular black hole geometry that avoids curvature singularities at the centre. The $(\beta+r)^{4}$ dependence ensures that the energy density remains finite as $r\to 0$ and falls off rapidly as $r\to\infty$, thereby satisfying the necessary energy conditions and preserving the asymptotic (A)dS structure.
@@ III Temperature
At the horizon $\mathcal{F}(r)=0\big{|}_{r=r_{h}}$, thus one can find the mass of the black hole as a function of the horizon radius.
@@
Hawking temperature is obtained from $\frac{1}{4\pi}\partial\mathcal{F}(r)/\partial r|_{r=r_{h}}$. For the metric of the form 6, by substituting the mass from Eq. (7), the temperature would be found as
$T_{H}={\frac{\beta^{4}+r_{h}^{4}+4\beta r_{h}^{3}-\alpha r_{h}^{2}+6\beta^{2}r_{h}^{2}-\Lambda r_{h}^{2}(\beta+r_{h})^{4}+4\beta^{3}r_{h}}{4\pi r_{h}(\beta+r_{h})^{4}}}$ (8)
@@
Fig. 3 illustrates the allowable range for the parameters $\alpha$ and $\beta$, for which there is a remnant radius for a specific value of parameter $\Lambda$.
Figure 3: Region where remnant radius does exist, assuming $\Lambda=0.01$.
By substituting the remnant radius obtained from Eq. (9), by $r_{h}$ in Eq. (7), the remnant mass of the black hole can be expressed as follows
@@ IV Photon and Shadow Radii
where the Lagrangian is obtained as Ahmed
@@
Thus, one can define an effective potential which is formulated as
@@
To ensure a circular orbit, two constraints are applied, which are as follows
@@
Thus, one can find that the following constraints should be satisfied
$E/L$ $=\pm\sqrt{\mathcal{F}(r)/r^{2}},$ (17)
@@
$r_{sh}=r_{ph}\frac{\sqrt{f(r_{o})}}{\sqrt{f(r_{ph})}},$ (19)
that $r_{o}$ indicates the observer radius. Shadow radius is shown in Figs. 4 and 5. Table 1, represents the permissible range of parameters based on the observation of Sgr A for $1\sigma$ and $2\sigma$ regions.
@@
Figure 4: Variation of shadow radius with parameter $\alpha$ for $r_{o}=10M$. In the left panel, beta=0.8\ beta=0.8 and in the right panel $\Lambda=-0.002$ are considered.
@@
As evident in Figs. 4 to 6, a reduction in the parameters $\alpha$, and $\Lambda$, while the rest of the parameters remain constant, results in a decrease in the value of the shadow radius.
[…]
@@
$\beta=0.900$ --- --- $0.720$ --- $\Lambda=-0.002$ --- --- $0.493$ ---
$\beta=0.030$ --- --- $0.508$ $0.813$ $\Lambda=-0.006$ --- --- $0.959$ ---
$\beta=0.010$ --- --- $0.415$ $0.653$ $\Lambda=-0.010$ --- $0.156$ --- ---
@@
$\alpha=0.900$ 0.041-0.041 0.030-0.030 0.014-0.014 0.007-0.007 $\beta=0.900$ 0.022-0.022 0.014-0.014 0.003-0.003 ---
$\alpha=0.500$ 0.025-0.025 0.017-0.017 0.005-0.005 --- $\beta=0.500$ 0.024-0.024 0.016-0.016 0.004-0.004 ---
$\alpha=0.010$ 0.015-0.015 0.009-0.009 --- --- $\beta=0.050$ 0.028-0.028 0.019-0.019 0.006-0.006 0.001-0.001
@@ V Massless Scalar and Electromagnetic Perturbations in Black Hole Spacetimes
so that, now, $l$ becomes the multipole moment of the black hole’s QNMs. The frequency $\omega$ in the stationary Schrödinger equation (21) is defined via $\partial_{t}^{2}\psi_{sl}(t,r)=-\omega^{2}\psi_{sl}(t,r)$ and this definition is made possible by the diagonal and static nature of the metric in (4). The wave function $\psi(r_{*})_{s}$ stands for the waves associated with the scalar QNMs Chandrasekhar:1985kt ; lopez2020 . The maximum of this potential is located at $\dfrac{\partial}{\partial r}\mathcal{V}_{s}(r)\bigg{|}_{r=r_{peak}}=0$.
@@
This effective potential $\mathcal{V}_{em}(r)$ of vector perturbations differs from the effective potential $\mathcal{V}_{s}(r)$ of scalar perturbations by the presence of the derivative of $g_{tt}$ in the latter, and the peak of the potential $\mathcal{V}_{em}(r)$ is located at $\dfrac{\partial}{\partial r}\mathcal{V}_{em}(r)\bigg{|}_{r=r_{peak}}=0$.
@@
In Fig.s 7, 8 and 9, we have depicted the behaviours of the scalar and electromagnetic potentials with respect to the model parameters $l$, $\alpha$ and $\beta$ respectively. From Fig. 7, we can see that the potentials associated with both scalar and electromagnetic potentials behave similarly. However, for the scalar perturbation, the value of the peak of the potential is higher than that of the electromagnetic potential.
@@
Table 2: The QNMs for the massless scalar and electromagnetic perturbation with $n=0$, $M=1$, $\alpha=0.3$, $\beta=0.2$ and $\Lambda=0.0$ using the Padé averaged 6th order WKB approximation method.
$l$ Scalar QNMs EM QNMs $\vartriangle_{rms}$(scalar) $\vartriangle_{rms}$(em) $\Delta_{6}$(scalar) $\Delta_{6}$(em)
1 0.3072010.0995284i0.307201-0.0995284i 0.2617420.0946083i0.261742-0.0946083i $1.13985\times 10^{-6}$ $7.76923\times 10^{-6}$ $3.51816\times 10^{-6}$ $6.88485\times 10^{-6}$
2 0.5071040.0986576i0.507104-0.0986576i 0.4805890.0969834i0.480589-0.0969834i $3.78281\times 10^{-7}$ $5.01194\times 10^{-7}$ $2.01920\times 10^{-6}$ $5.55769\times 10^{-7}$
3 0.7080880.0984074i0.708088-0.0984074i 0.6892860.0975634i0.689286-0.0975634i $38391\times 10^{-8} 8.8$$84296\times 10^{-8}$ $3.96363\times 10^{-7}$ $1.05988\times 10^{-7}$
4 0.9094200.0983032i0.909420-0.0983032i 0.8948390.0977950i0.894839-0.0977950i $2.42110\times 10^{-8}$ $2.46254\times 10^{-8}$ $1.13709\times 10^{-7}$ $3.05768\times 10^{-8}$
5 1.1109100.0982502i1.110910-0.0982502i 1.0990000.0979107i1.099000-0.0979107i $8.87193\times 10^{-9}$ $8.92710\times 10^{-9}$ $4.17001\times 10^{-8}$ $1.12966\times 10^{-8}$
6 1.3124800.0982197i1.312480-0.0982197i 1.3024100.0979769i1.302410-0.0979769i $3873\times 10^{-9} 3.838$$3890\times 10^{-9}$ $4.90337\times 10^{-9}$ $4.92176\times 10^{-9}$
7 1.5141100.0982005i1.514110-0.0982005i 1.5053800.0980182i1.505380-0.0980182i $3.12285\times 10^{-9}$ $3.12212\times 10^{-9}$ $2.40862\times 10^{-9}$ $2.41337\times 10^{-9}$
8 1.7157600.0981876i1.715760-0.0981876i 1.7080700.0980458i1.708070-0.0980458i $1.47073\times 10^{-9}$ $1.47008\times 10^{-9}$ $1.29188\times 10^{-9}$ $1.29362\times 10^{-9}$
@@
We have listed QNMs for $n=0$ and for different $l$ values in Table 2 for both scalar and electromagnetic perturbations. These results were obtained for deformation parameters $\alpha=0.3$ and $\beta=0.2$, which influence the spacetime geometry and black hole properties. The table displays QNM frequencies, with the real part representing the oscillation frequency and the imaginary part corresponding to the damping rate of the modes. These modes were calculated for different multipole numbers $l$, ranging from $l=1$ to $l=8$. To ensure the reliability of our analysis, we have carefully chosen parameter values that maintain a well-defined, single-peak potential structure, which is a prerequisite for the WKB method to yield meaningful results.
@@
As expected, the real part of the QNMs increases with $l$, indicating higher oscillation frequencies for modes with larger angular momentum. The imaginary part, which controls the damping or decay rate, shows slight variation across the different values of $l$. Interestingly, scalar perturbations consistently exhibit slightly higher oscillation frequencies and more pronounced damping compared to electromagnetic perturbations. This is reflected in the larger imaginary components for scalar modes compared to their electromagnetic counterparts, highlighting the different behaviours of these two types of perturbations in deformed black hole spacetimes.
@@
The table also presents the root mean square (rms) errors, $\Delta_{rms}$, for both scalar and electromagnetic perturbations. These errors quantify the precision of the WKB method in estimating the QNMs. It is observed that the RMS errors generally decrease with increasing $l$, indicating more accurate QNM calculations for higher multipole numbers. The WKB errors, $\Delta_{6}$, which are basically an error term associated with half of the absolute difference between QNMs calculated using 5th and 7th order WKB methods, similarly decrease with increasing $l$. This suggests that the WKB method becomes more reliable for larger values of $l$, where the approximation yields more precise QNM frequencies. Nevertheless, the electromagnetic perturbations exhibit slightly larger errors compared to scalar perturbations, although both remain within acceptable limits, particularly for $l\geq 2$.
@@
In Fig. 10, we have shown the variation of QNMs with respect to the model parameter $\alpha$ for scalar perturbations. On the first panel, we have shown the variations of the real part of QNMs i.e. the oscillation frequencies of ring-down GWs with respect to the model parameter $\alpha$. With an increase in the value of $\alpha$, oscillation frequency increases significantly. In the case of imaginary QNMs, we can see that an increase in the value of $\alpha$ increases the damping rate of ring-down GWs. In Fig. 11, we have depicted the variations of QNMs with respect to the model parameter $\beta$. On the first panel of Fig. 11, one can see that with an increase in the value of $\beta$, the real part of QNMs, i.e. oscillation frequency of GWs decreases non-linearly. In the case of damping rate, we observe an interesting behaviour. Initially, the damping rate of the decay rate of ring-down GWs increases with an increase in the value of $\beta$ up to around $\beta=0.65.$ Beyond this value, the decay rate decreases with an increase in the model parameter $\beta$. In the cases of electromagnetic perturbations, we observe a similar scenario with slightly lower values of oscillation frequencies as shown in Fig.s 12 and 13. This feature arises due to the dual role of $\beta$, which not only ensures regularity at the core by avoiding the central singularity but also modifies the effective potential governing scalar perturbations. Initially, increasing $\beta$ enhances the damping (i.e., $\omega_{I}$ becomes more negative) due to a steeper potential barrier; however, beyond a critical value, the core regularization becomes dominant, leading to a broadening or softening of the potential that supports longer-lived modes, hence a reversal in the damping trend. This behaviour may signal a transition in the internal structure of the spacetime and could carry potential astrophysical implications if such geometries model exotic compact objects—offering observable signatures in gravitational wave ringdowns that deviate from classical black hole predictions.
@@
Now, we would like to investigate the time domain profiles of the perturbations. To do so, we shall use the scheme introduced in Ref. gundlach . This methodology allows us to investigate the temporal evolution and decay characteristics of perturbations, particularly at late times, thereby furnishing an independent check on the stability and consistency of the system under consideration.
To implement this technique, we discretize the wave equation on a uniform grid by defining the scalar field as $\psi(r_{*},t)=\psi(i\Delta r_{*},j\Delta t)=\psi_{i,j}$, where $i$ and $j$ index the spatial and temporal grid points respectively. The effective potential is similarly discretized as $V(r(r_{*}))=V(r_{*},t)=V_{i,j}$, which in the time-independent case simplifies to $V_{i,j}=V_{i}$.
Using these discretized variables, the scalar wave equation takes the following finite difference form:
$\frac{\psi_{i+1,j}-2\psi_{i,j}+\psi_{i-1,j}}{\Delta r_{*}^{2}}-\frac{\psi_{i,j+1}-2\psi_{i,j}+\psi_{i,j-1}}{\Delta t^{2}}-V_{i}\psi_{i,j}=0.$ (27)
This equation can be explicitly solved for $\psi_{i,j+1}$, yielding the following iterative scheme for the time evolution of the perturbation:
$\psi_{i,j+1}=-\psi_{i,j-1}+\left(\frac{\Delta t}{\Delta r_{*}}\right)^{2}(\psi_{i+1,j}+\psi_{i-1,j})+\left[2-2\left(\frac{\Delta t}{\Delta r_{*}}\right)^{2}-V_{i}\Delta t^{2}\right]\psi_{i,j}.$ (28)
The initial profile for the scalar perturbation is modelled as a Gaussian wave packet:
$\psi(r_{*},t=0)=\exp\left[-\frac{(r_{*}-k_{1})^{2}}{2\sigma^{2}}\right],$ (29)
with the condition $\psi(r_{*},t<0)=0$, where $k_{1}$ and $\sigma$ denote the central location and width of the initial packet, respectively.
To ensure numerical stability of the iterative process, the ratio $\Delta t/\Delta r_{*}$ must satisfy the Von Neumann stability criterion, i.e.,
$\frac{\Delta t}{\Delta r_{*}}<1.$ (30)
This constraint ensures convergence of the solution and prevents the amplification of numerical errors during evolution.
By employing this finite-difference scheme, we numerically evolve the perturbation profile over time, thereby obtaining the time-domain signal. This evolution not only facilitates the identification of quasi-normal ringing but also serves as a crucial cross-validation for the quasi-normal mode spectrum derived in the frequency domain. Consequently, the combined analysis offers a more robust and comprehensive understanding of the perturbative dynamics in the considered gravitational background.
Refer to caption    Refer to caption
Figure 14: The time-domain profiles of the massless scalar perturbations (first panel) and electromagnetic perturbations (second panel) for different multipole moments $l$ with the parameter values $M=1$, $G=1$, $n=0$, $\alpha=0.3$, $\beta=0.1$ and $\Lambda=-0.002$.
Refer to caption    Refer to caption
Figure 15: The time-domain profiles of the massless scalar perturbations (first panel) and electromagnetic perturbations (second panel) for different values of $\alpha$ with the parameter values $M=1$, $G=1$, $n=0$, $l=1$, $\beta=0.1$ and $\Lambda=-0.002$.
Refer to caption    Refer to caption
Figure 16: The time-domain profiles of the massless scalar perturbations (first panel) and electromagnetic perturbations (second panel) for different values of $\beta$ with the parameter values $M=1$, $G=1$, $n=0$, $\alpha=0.3$, $l=1$ and $\Lambda=-0.002$.
Refer to caption    Refer to caption
Figure 17: The time-domain profiles of the massless scalar perturbations (first panel) and electromagnetic perturbations (second panel) for different multipole moments $l$ with the parameter values $M=1$, $G=1$, $n=0$, $\alpha=0.3$, $\beta=0.1$ and $\Lambda=0.0$.
Refer to caption    Refer to caption
Figure 18: The time-domain profiles of the massless scalar perturbations (first panel) and electromagnetic perturbations (second panel) for different values of $\alpha$ with the parameter values $M=1$, $G=1$, $n=0$, $l=1$, $\beta=0.1$ and $\Lambda=0.0$.
Refer to caption    Refer to caption
Figure 19: The time-domain profiles of the massless scalar perturbations (first panel) and electromagnetic perturbations (second panel) for different values of $\beta$ with the parameter values $M=1$, $G=1$, $n=0$, $\alpha=0.3$, $l=1$ and $\Lambda=0.0$.
We have shown the time domain profiles for scalar and electromagnetic perturbations in Fig.s 14, 15, 16, 17, 18 and 19. The time-domain profiles presented in these figures exhibit strong consistency with the QNM frequencies computed using the sixth-order WKB-Padé approximation, as listed in Table 2. For fixed deformation parameters $\alpha=0.3$, $\beta=0.2$, and $\Lambda=0$, the scalar and electromagnetic perturbations display oscillation frequencies and damping rates in the time-domain evolution that closely match the real and imaginary parts of the fundamental ($n=0$) QNMs. As the multipole index $l$ increases, the time-domain signals show higher oscillation frequencies and slower damping, which aligns well with the increasing real part and decreasing imaginary part of the QNM spectrum. Scalar perturbations consistently exhibit slightly higher frequencies and faster decay than electromagnetic ones, a trend that is evident in both the numerical evolution and the WKB results. Additionally, the influence of the deformation parameter $\alpha$ and the control parameter $\beta$ is reflected in the time-domain profiles: increasing $\alpha$ leads to more pronounced peaks and slightly delayed decay, while increasing $\beta$ tends to reduce the oscillation frequency and prolong the signal, indicating a flatter effective potential. These qualitative behaviors are in agreement with theoretical expectations from perturbation theory, even though the WKB data is only presented for fixed parameter values. Overall, the time-domain analysis not only corroborates the WKB-based QNM frequencies but also provides deeper insight into the dynamical response of deformed Schwarzschild black holes, reinforcing the reliability of both methods in capturing the essential features of scalar and electromagnetic field perturbations.